The smaller of two numbers is 2 less than 1/3 the larger. The sum. of the two numbers is 50. Find the value of the numbers.
Answer by TimothyLamb(4379)

s = (1/3)b - 2
s + b = 50
s = 50 - b
---
(1/3)b - 2 = 50 - b
(1/3)b - 2 = 50 - (3/3)b
(4/3)b = 52
b = 52(3/4)
---
b = 39
s = 11
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php